# Flow-based Conformal Prediction for Multi-dimensional Time Series

### v2

@@ Abstract
Time series prediction underpins a broad range of downstream tasks across many scientific domains. Recent advances and increasing adoption of black-box machine learning models for time series prediction highlight the critical need for reliable uncertainty quantification. While conformal prediction has gained attention as a reliable uncertainty quantification method, conformal prediction for time series faces two key challenges: (1) adaptively leveraging correlations in features and non-conformity scores to overcome the exchangeability assumption, and (2) constructing prediction sets for multi-dimensional outcomes. To address these challenges jointly, we propose a novel conformal prediction method for time series using flow with classifier-free guidance. We provide coverage guarantees by establishing exact non-asymptotic marginal coverage and a finite-sample bound on conditional coverage for the proposed method. Evaluations on real-world time series datasets demonstrate that our method constructs significantly smaller prediction sets than existing conformal prediction methods while maintaining target coverage.
@@ 1 Introduction
Figure 1: Our method adaptively constructs the prediction set at time $i$ using a flow transformation $\psi$ conditioned on guidance $h_{i}$, which encodes contextual information extracted from past features and residuals.
@@
Uncertainty quantification has become essential in scientific fields where black-box machine learning models are widely deployed [angelopoulos2021gentle]. Conformal prediction (CP) has emerged as a reliable, distribution-free framework for uncertainty quantification that constructs prediction sets with coverage guarantees, ensuring they contain the true outcome with a specified confidence level [shafer2008tutorial, vovk2005algorithmic]. By constructing uncertainty sets using non-conformity scores that quantify how atypical predictions are, CP generates reliable prediction sets that satisfy a specified confidence level.
@@
Time series prediction aims to forecast future outcomes based on past sequential observations of features [box2015time], and underpins a broad range of downstream tasks. Recent advances in machine learning have led to the development of various foundation models designed for time series prediction [kim2025comprehensive, miller2024survey, wen2023transformers]. The growing adoption of such models for time series prediction highlights the pressing need for reliable uncertainty quantification. Although CP has been actively studied for reliable uncertainty quantification, most existing CP methods rely on the assumption of data exchangeability. The exchangeability assumption is frequently violated in time series data, where observations exhibit complex temporal dependencies that induce correlations in the non-conformity scores, thereby making the direct application of CP to time series prediction particularly challenging. An additional challenge is that modern time series data often contain high-dimensional features and multi-dimensional outcomes. While CP methods for univariate outcomes are well-established, extending these methods to generate prediction sets for multi-dimensional outcomes is not straightforward and requires careful consideration in constructing prediction sets.
@@
There has been substantial effort to extend CP beyond the exchangeability assumption. One line of research focuses on addressing distribution shifts in the data [barber2023conformal, tibshirani2019conformal]. More recently, several works have developed CP methods for time series. For example, xu2021conformal proposed a method to construct sequential prediction intervals for time series based on a bootstrap ensemble estimator, which were later extended to incorporate conditional quantile estimation in order to exploit correlations in non-conformity scores [xu2023sequential]. auer2024conformal used modern Hopfield networks to capture temporal dependencies by reweighting samples, and constructed prediction intervals based on these reweighting. Another line of work have proposed multi-step conformal prediction methods for time series, but they assume access to multiple i.i.d. sequences of time series [stankeviciute2021conformal, sun2022copula], which may limit their applicability in general practical settings.
@@
Constructing prediction sets for multi-dimensional outcomes has been an active area of research. Early approaches used copulas [messoudi2021copula] and ellipsoidal uncertainty sets [henderson2024adaptive, johnstone2022exact, messoudi2022ellipsoidal], yielding hyper-rectangular and ellipsoidal prediction sets, respectively. Subsequent research has aimed to move beyond specific geometric shapes of prediction sets: braun2025minimum formulated structured non-convex optimization to obtain minimum-volume sets; and tumu2024multi used convex templates for prediction sets. Recent works have focused on transporting multi-dimensional non-conformity scores to a reference distribution from which prediction sets can be constructed. For example, klein2025multivariate and thurinoptimal used Monge–Kantorovich ranks [chernozhukov2017monge, hallin2021distribution] to map multi-dimensional non-conformity scores onto a reference distribution to construct prediction sets, by solving optimal transport problems. fang2025contra applied conditional normalizing flows to map multi-dimensional non-conformity scores to the source distribution and construct prediction sets using a calibration set with the source distribution.
@@
Consequently, an effective CP method for time series prediction must address the two aforementioned challenges simultaneously: leveraging correlations in features and non-conformity scores to overcome the exchangeability assumption, and constructing prediction sets for multi-dimensional outcomes. To the best of our knowledge, xu2024conformal is the only work that seeks to address both challenges jointly, constructing ellipsoidal prediction sets by defining non-conformity scores as the radii of ellipsoidal sets and predicting these non-conformity scores conditionally.
@@
In this work, we propose a novel conformal prediction method designed for time series prediction with multi-dimensional outcomes. Our method is designed to effectively address the aforementioned two challenges by using flow with classifier-free guidance. Specifically, we use flow to model the distribution of prediction residuals and their transformations conditioned on historical context, which is encoded by using Transformer. We define the non-conformity score as the Euclidean distance between the transformed prediction residual and the mean of a Gaussian source distribution of the flow, which allows us to construct prediction sets at a desired confidence level. We provide theoretical coverage guarantees by establishing exact non-asymptotic marginal coverage and a finite-sample bound on conditional coverage for the proposed method. Empirical evaluations on three real-world multi-dimensional time series datasets demonstrate that the proposed method constructs significantly smaller prediction sets while maintaining target coverage, outperforming existing baselines.
@@
Figure 2: Comparison of the prediction sets with a target coverage of 0.95, constructed by FCP (ours), MultiDimSPCI [xu2024conformal], and conformal prediction using copula [messoudi2021copula] on (a) wind, (b) traffic, and (c) solar datasets. Prediction sets are manually selected from the test set for visual clarity. Two prediction sets are shown for the wind dataset.
@@ 2 Problem Setup
Suppose that the first $T$ examples, $\{(x_{i},y_{i})\}_{i=1}^{T}$, are used for training. Our goal is to sequentially construct a prediction set $\hat{C}_{i}(z_{i},\alpha)$ for the next step, beginning at time $i=T+1$. Here, $z_{i}$ denotes the features used to construct $\hat{C}_{i}$, and $\alpha\in[0,1]$ denotes a pre-specified significance level. In the simplest setting, $z_{i}$ consists only of $x_{i}$, but it may also include past features or non-conformity scores. We aim to construct prediction sets that satisfy marginal coverage:
@@
Although trivially large prediction sets can always satisfy the target coverage, they do not provide useful information for uncertainty quantification. Therefore, the meaningful objective is to construct efficient prediction sets—the prediction sets that are as small as possible while satisfying the target coverage [vovk2005algorithmic].
@@ 3.1 Preliminary: Guided Flow
We use $x$ as a generic variable in this section, distinct from the time series feature $x_{i}$ introduced in the problem setup. A flow is a time-dependent mapping $\psi:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that push-forward a random variable $X_{0}\in\mathbb{R}^{d}$ from a source distribution $p_{0}$ to $X_{t}\in\mathbb{R}^{d}$ from a time-dependent probability density (i.e., probability path) $p_{t}$ for time $t\in[0,1]$ as follows:
@@
One can verify the marginal guided vector field generates the marginal guided probability path using the continuity equation (see Proposition A.1). Therefore, in order to learn the target conditional distribution, we parameterize the guided vector field with neural networks and train it to approximate the marginal guided vector field as closely as possible. A simple and effective way to train the guided vector field is through flow matching, which minimizes the mean-squared error between the conditional guided vector field and the parameterized guided vector field [lipman2022flow]:
@@
where $t\sim\text{Unif}[0,1]$, $(x_{1},h)\sim q_{\text{data}}$, and $u_{t|h}^{\theta}$ is the parameterized guided vector field with parameters $\theta$.
@@
where $u_{t}(x)$ is unguided vector field, $b_{t}$ is a scalar constant regarding $\alpha_{t}$ and $\sigma_{t}$ (see Proposition 36). Based on this reformulation, early approaches trained a separate classifier [song2020score] with a classifier scale $w>1$ is beneficial in conditional generation in practice [dhariwal2021diffusion] :
@@
Instead of modeling $u_{t}(x)$ and $u_{t}(x\mid h)$ separately, ho2022classifier proposed using a single vector field to model both cases by assigning a null condition $h_{\varnothing}$ to represent the unguided vector field, which is known as classifier-free guidance (CFG):
@@
where $\eta\sim\text{Bernoulli}(p_{\varnothing})$ and $p_{\varnothing}$ denotes the probability of assigning $h_{\varnothing}$. The resulting guided vector field $\tilde{u}_{t|h}(x\mid h)$ in equation (12) enables conditional generation by solving the guided flow ODE and has been widely used in various tasks such as image generation [esser2024scaling] and video generation [polyak2025moviegencastmedia].
3.2 Conformal Prediction for Time Series using Guided Flow
We use guided flow to learn a mapping from the source distribution to the distribution of prediction residual $\hat{\epsilon}=y-\hat{y}$, conditioned on past features and residuals. The prediction set is then defined through this transformation using guided flow to achieve the target coverage. This construction effectively addresses the two aforementioned key challenges in conformal prediction for time series. First, the guided flow explicitly captures correlations among past features and residuals by using them as guidance. Second, since the transformation using the guided flow can be defined between random variables in arbitrary dimensions, it enables the generation of prediction sets for multi-dimensional outcomes in any $\mathbb{R}^{d_{y}}$. Figure 1 provides a visual illustration of the method. We describe the method in detail in this section.
Guided flow design.
We use Gaussian probability path with interpolating scheduler $a_{t}=t$ and $\sigma_{t}=(1-t)$. The source distribution is set to an isotropic Gaussian with zero mean and covariance scale $\gamma>0$, i.e., $\mathcal{N}(0,\gamma I_{d_{y}})$. For each time index $i$, we construct $z_{i}$ by concatenating the past $w$ features and prediction residuals, and use an encoder to obtain a contextual representation $h_{i}=\text{Enc}(z_{i})$. The classifier-free guided vector field as defined in equation (12) uses $h_{i}$ as the guidance to model the conditional distribution of $\hat{\epsilon}_{i}$. In our method, we use Transformer as the encoder [vaswani2017attention], though alternative sequence models such as recurrent neural networks (RNNs) are also applicable. The guided vector field is trained via flow matching as defined in equation (13), and the encoder is jointly trained with it. The overall training procedure is summarized in Algorithm 1.
@@ 3.2 Conformal Prediction for Time Series using Guided Flow
 $h_{i}\leftarrow\text{Enc}^{\theta}(z_{i})$ $\triangleright$ obtain contextual representation
@@
The trained guided flow models the conditional distribution of the prediction residual by mapping samples from the source Gaussian distribution to residuals conditioned on the guidance $h_{i}$. Since this transformation is bijective, we can define prediction sets for the residuals directly through the transformation. Let $\hat{e}_{i}(y):=||\psi^{-1}_{t=1|h}(\hat{\epsilon}\mid h_{i})||$ be the Euclidean distance between the transformed residual and the origin, then the prediction set at significance level $\alpha$ can be defined as:
@@
Since the prediction set is obtained through the transformation using the guided flow, it can take on flexible shapes without being constrained to follow any fixed geometric form, such as convex or ellipsoidal sets. We believe this enables the guided flow to generate smaller prediction sets that are better aligned with the data and the guidance. Although the prediction sets do not have any fixed geometric shape, some useful topological properties can still be inferred. In particular, Theorem A.4 ensures that the prediction sets are closed and connected. Figure 2 shows prediction sets in $\mathbb{R}^{2}$ generated by our proposed method alongside two other methods that produce hyper-rectangular prediction sets [messoudi2021copula] and ellipsoidal prediction sets [xu2024conformal]. The figure visually demonstrates the flexible shapes of the prediction sets constructed by our proposed method.
@@
where $\psi_{1}$ represents the flow transformation from $t=0$ to $t=1$, and $J_{\psi_{1}}(x\mid h)$ denotes the Jacobian of $\psi_{1}$ at $x\in\mathcal{B}_{\alpha}$ conditioned on $h$. The right-hand side provides a Monte Carlo approximation, where $x_{j}$ are i.i.d. samples drawn from $\mathcal{B}_{\alpha}$ and $N$ is the number of samples. However, directly computing $\det\left(J_{\psi_{1}}(x\mid h)\right)$ is computationally expensive, as it requires solving the guided flow ODE and evaluating the full Jacobian matrix. Instead, we can compute the log-determinant of the Jacobian by solving the following ODE:
@@
where $\textrm{div}(\cdot)$ denotes the divergence operator. A detailed derivation is provided in Proposition A.3. The accuracy of the prediction set size estimate depends on the Monte Carlo approximation. Purely random sampling from $\mathcal{B}_{\alpha}$ may introduce bias due to uneven coverage of the sampling space, and a small sample size $N$ can result in high variance. To reduce sampling bias, we use quasi-Monte Carlo sampling based on Sobol sequences [sobol1967distribution, practicalqmc], which provides more uniform sampling from $\mathcal{B}_{\alpha}$. To control variance from finite sampling, we monitor the relative error in terms of sample size $N$. Additional implementation details are provided in the experiment section.

### v3

@@ Abstract
Time series prediction underpins a broad range of downstream tasks across many scientific domains. Recent advances and increasing adoption of black-box machine learning models for time series prediction highlight the critical need for uncertainty quantification. While conformal prediction has gained attention as a reliable uncertainty quantification method, conformal prediction for time series faces two key challenges: (1) leveraging correlations in observations and non-conformity scores to overcome the exchangeability assumption, and (2) constructing prediction sets for multi-dimensional outcomes. To address these challenges, we propose a novel conformal prediction method for time series using flow with classifier-free guidance. We provide coverage guarantees by establishing exact non-asymptotic marginal coverage and a finite-sample bound on conditional coverage for the proposed method. Evaluations on real-world time series datasets demonstrate that our method constructs significantly smaller prediction sets than existing conformal prediction methods, maintaining target coverage.
@@ 1 Introduction
Figure 1: Our method adaptively constructs the prediction set using a flow transformation $\psi$ conditioned on guidance $h_{i}$, which encodes the past context.
@@
Uncertainty quantification has become essential in scientific fields where black-box machine learning models are widely deployed [1]. Conformal prediction (CP) has emerged as a distribution-free framework for uncertainty quantification with coverage guarantees, ensuring prediction sets contain the true outcome at a specified confidence level [39, 48]. By constructing uncertainty sets using non-conformity scores that quantify how atypical predictions are, CP generates reliable prediction sets that satisfy a specified confidence level.
@@
Time series prediction aims to forecast future outcomes from past sequential observations of features and outcomes [7], underpinning a broad range of downstream tasks. Recent advances in machine learning have led to the development of various foundation models designed for time series prediction [24, 32, 49]. The growing adoption of such models for time series prediction highlights the pressing need for reliable uncertainty quantification. Although CP has been actively studied for reliable uncertainty quantification, most existing CP methods rely on the assumption of data exchangeability [3]. The exchangeability assumption is frequently violated in time series data, where observations exhibit complex temporal dependencies that induce correlations in the non-conformity scores, thereby making the direct application of CP to time series prediction particularly challenging. An additional challenge is that modern time series data often contain multi-dimensional outcomes. While CP methods for univariate outcomes are well-established, extending these methods to generate prediction sets for multi-dimensional outcomes is not straightforward and requires careful consideration in constructing prediction sets.
@@
There has been substantial effort to extend CP beyond the exchangeability assumption. One line of research focuses on addressing distribution shifts in the data [3, 45]. More recently, several works have developed CP methods for time series. For example, Xu and Xie [52] proposed a method to construct sequential prediction intervals for time series based on a bootstrap ensemble estimator, which were later extended to incorporate conditional quantile estimation in order to exploit correlations in non-conformity scores [54]. Auer et al. [2] used modern Hopfield networks to capture temporal dependencies for sample reweighting and constructed prediction intervals based on these weights. Another line of work has proposed multi-step conformal prediction methods for time series, but these approaches assume access to multiple i.i.d. time series sequences [42, 43], which may limit their applicability in practical settings. Despite these efforts, most of the existing methods remain limited to univariate outcomes.
@@
Constructing prediction sets for multi-dimensional outcomes has been an active area of research. Early approaches used copulas [30] and ellipsoidal uncertainty sets [19, 23, 31], yielding hyper-rectangular and ellipsoidal prediction sets, respectively. Subsequent research has aimed to move beyond specific geometric shapes of prediction sets: Braun et al. [8] formulated structured non-convex optimization to obtain minimum-volume sets; and Tumu et al. [46] used convex templates for prediction sets. Recent works have focused on transporting multi-dimensional non-conformity scores to a reference distribution from which prediction sets can be constructed. For example, Klein et al. [26] and Thurin et al. [44] used Monge–Kantorovich ranks [11, 18] to map multi-dimensional non-conformity scores onto a reference distribution to construct prediction sets, by solving optimal transport problems. Fang et al. [16] applied conditional normalizing flows to map multi-dimensional non-conformity scores to the source distribution and construct prediction sets using a calibration set with the source distribution.
@@
Consequently, an effective CP method for time series prediction should address the two aforementioned challenges: overcoming the exchangeability assumption and constructing prediction sets for multi-dimensional outcomes. To the best of our knowledge, Xu et al. [50] is the only work that seeks to jointly addresses both challenges, constructing ellipsoidal prediction sets by defining non-conformity scores as the radii of the sets and predicting these non-conformity scores conditionally.
@@
In this work, we propose a novel conformal prediction method for time series prediction with multi-dimensional outcomes. Our method is designed to effectively address the two challenges by using flow with classifier-free guidance. Specifically, we use flow to model the distribution of prediction residuals and their transformations conditioned on past context, which is encoded by using Transformer [47]. We define the non-conformity score as the Euclidean distance between the transformed prediction residual and the mean of a Gaussian source distribution of the flow, which allows us to construct prediction sets at a desired confidence level. We provide theoretical coverage guarantees by establishing an exact non-asymptotic marginal coverage and a finite-sample bound on conditional coverage for the proposed method. Empirical evaluations on three real-world multi-dimensional time series datasets demonstrate that the proposed method constructs significantly smaller prediction sets while maintaining target coverage, outperforming existing baselines.
@@
Figure 2: Comparison of the prediction sets at a target coverage of 0.95, constructed by FCP (ours), MultiDimSPCI [50], and conformal prediction using empirical copula [30] on (a) wind, (b) traffic, and (c) solar datasets. Prediction sets are manually selected from the test set for visual clarity. Two prediction sets are shown for the wind dataset.
@@ 2 Problem Setup
Suppose that the first $T$ examples, $\{(x_{i},y_{i})\}_{i=1}^{T}$, are used for training. Our goal is to sequentially construct a prediction set $\hat{C}_{i}(z_{i},\alpha)$ for the next step, beginning at time $i=T+1$. Here, $z_{i}$ denotes the features used to construct $\hat{C}_{i}$, and $\alpha\in[0,1]$ denotes a pre-specified significance level. In the simplest setting, $z_{i}$ consists only of $x_{i}$, but it may also include past features, outcomes, or non-conformity scores. We aim to construct prediction sets that satisfy marginal coverage:
@@
Although trivially large prediction sets can always satisfy the target coverage, they do not provide useful information for uncertainty quantification. Therefore, the meaningful objective is to construct efficient prediction sets—the prediction sets that are as small as possible while satisfying the target coverage [48].
@@ 3.1 Preliminary: Guided Flow
We use $x$ as a generic variable in this section, distinct from the time series feature $x_{i}$ introduced in the problem setup. A flow is a time-dependent mapping $\psi:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ that push-forward a random variable $X_{0}\in\mathbb{R}^{d}$ from a source distribution $p_{0}$ to $X_{t}\in\mathbb{R}^{d}$ from a time-dependent probability density (i.e., probability path) $p_{t}$ at time $t\in[0,1]$ as follows:
@@
One can verify the marginal guided vector field generates the marginal guided probability path using the continuity equation (see Proposition A.1). Therefore, in order to learn the target conditional distribution, we parameterize the guided vector field with neural networks and train it to approximate the marginal guided vector field as closely as possible. A simple and effective way to train the guided vector field is through flow matching, which minimizes the mean-squared error between the conditional guided vector field and the parameterized guided vector field [29]:
@@
where $t\sim\text{Unif}[0,1]$, $(x_{1},h)\sim q_{\text{data}}$, and $u_{t|h}^{\theta}$ is the guided vector field with parameters $\theta$.
@@
where $u_{t}(x)$ is unguided vector field, $b_{t}$ is a scalar constant regarding $\alpha_{t}$ and $\sigma_{t}$ (see Proposition 32). Based on this reformulation, early approaches trained a separate classifier [41] with a classifier scale $w>1$ is beneficial in conditional generation in practice [12] :
@@
Instead of modeling $u_{t}(x)$ and $u_{t}(x\mid h)$ separately, Ho and Salimans [21] proposed using a single vector field to model both cases by assigning a null condition $h_{\varnothing}$ to represent the unguided vector field, which is known as classifier-free guidance (CFG):
@@
where $\eta\sim\text{Bernoulli}(p_{\varnothing})$ and $p_{\varnothing}$ denotes the probability of assigning $h_{\varnothing}$. The resulting guided vector field $\tilde{u}_{t|h}(x\mid h)$ in equation (12) enables conditional generation by solving the guided flow ODE and has been widely used in various tasks such as image generation [15] and video generation [35].
@@
 $h_{i}\leftarrow\text{Enc}^{\theta}(z_{i})$ $\triangleright$ obtain guidance
@@
3.2 Conformal Prediction for Time Series using Guided Flow
We use guided flow to learn a mapping from the source distribution to the distribution of prediction residuals $\hat{\epsilon}=y-\hat{y}$, conditioned on past context. The prediction set is then defined through this transformation using guided flow to achieve the target coverage. This construction effectively addresses the two aforementioned challenges in conformal prediction for time series. First, guided flow leverages guidance from the encoder derived from past context, enabling accurate estimation of the conditional distribution of $\hat{\epsilon}$ and mitigating the reliance on the exchangeability assumption. Second, since guided flow defines transformations between random variables in arbitrary dimensions, it enables the generation of prediction sets for any multi-dimensional outcomes. Figure 1 provides a visual illustration of the method. We describe the method in detail in this section.
Guided flow design.
We use Gaussian probability path with interpolating scheduler $a_{t}=t$ and $\sigma_{t}=(1-t)$. The source distribution is set to an isotropic Gaussian with zero mean and covariance scale $\gamma>0$, i.e., $\mathcal{N}(0,\gamma I_{d_{y}})$. For each time index $i$, we construct $z_{i}$ by concatenating the past $w$ features and prediction residuals, and use an encoder to obtain a contextual representation $h_{i}=\text{Enc}(z_{i})$. The guided vector field as defined in equation (12) uses $h_{i}$ as the guidance to model the conditional distribution of $\hat{\epsilon}_{i}$. In our method, we use Transformer as the encoder [47], though alternative sequence models such as recurrent neural networks (RNNs) are also applicable. The guided vector field is trained via flow matching as defined in equation (13), and the encoder is jointly trained with it. The overall training procedure is summarized in Algorithm 1.
@@ 3.2 Conformal Prediction for Time Series using Guided Flow
The trained guided flow models the conditional distribution of the prediction residual by mapping from the source Gaussian distribution to residuals conditioned on the guidance. Since this transformation is bijective, we can define prediction sets for the residuals directly through the transformation. Let $\hat{e}_{i}(y):=||\psi^{-1}_{t=1|h}(\hat{\epsilon}\mid h_{i})||$ be the Euclidean distance between the transformed residual and the origin, then the prediction set at significance level $\alpha$ can be defined as:
@@
Since the prediction set is obtained through the transformation using guided flow, it can take on flexible shapes without being constrained to follow any fixed geometric form, such as convex or ellipsoidal sets. We believe this enables the guided flow to generate smaller prediction sets that are better aligned with the data. Although the prediction sets do not have any fixed geometric shape, some useful topological properties can still be inferred. In particular, Theorem A.4 and A.5 ensure that the prediction sets are closed and connected. Figure 2 shows prediction sets in $\mathbb{R}^{2}$ generated by our proposed method alongside two other methods that produce hyper-rectangular prediction sets [30] and ellipsoidal prediction sets [50]. The figure visually demonstrates the flexible shapes of the prediction sets constructed by our proposed method.
@@
where $\psi_{1}$ represents the flow transformation from $t=0$ to $t=1$, and $J_{\psi_{1}}(x\mid h)$ denotes the Jacobian of $\psi_{1}$ at $x\in\mathcal{B}_{\alpha}$ conditioned on $h$. The right-hand side provides a Monte Carlo approximation, where $x_{j}$ are i.i.d. samples drawn from $\mathcal{B}_{\alpha}$ and $N$ is the number of samples. However, directly computing $\det\left(J_{\psi_{1}}(x\mid h)\right)$ is computationally expensive, as it requires solving the guided flow ODE and evaluating the full Jacobian. Instead, we compute the log-determinant of the Jacobian by solving the following ODE:
@@
where $\textrm{div}(\cdot)$ denotes the divergence operator. A detailed derivation is provided in Proposition A.3. The accuracy of the prediction set size estimate depends on the Monte Carlo approximation. Purely random sampling from $\mathcal{B}_{\alpha}$ may introduce bias due to uneven coverage of the sampling space, and a small sample size $N$ can result in high variance. To reduce sampling bias, we use quasi-Monte Carlo sampling based on Sobol sequences [40, 34], which provides more uniform sampling from $\mathcal{B}_{\alpha}$. To control variance from finite sampling, we monitor the relative error in terms of sample size $N$. Additional implementation details are provided in the experiments section.